OpenStudy (anonymous):

a car is moving on a straight road with uniform acceleration .the following table gives the speed of the car at different time interval .time(sec)-0,10,20,30,40,50,60.speed-10,20,30,40,50,60,70.(a) plot a speed - time graph.(b) calculate the distance travelled by the car in 60 sec.(c) what do you infer from the graph.

OpenStudy (anonymous):

 time is on x-axis and speed on y-axis as the motion is 1D so distance travelled is area under the curve from x=0 to x=60 ie area of trapazeium=1/2(sum of parllel sides)*(perpendicular distance between them) =1/2(10+70)(60)=40*60=2400m from graph we infer that the acceleration is linear and away from the obsever:)